Iron can be obtained by reacting the ore hematite (Fe2O3) with coke (C).
The latter is converted to CO2. As manager of a blast furnace you are told that you have
20.5 Mg (megagrams) of Fe2O3 and 2.84 Mg of coke on hand. (a) Which should you order
first—another shipment of iron ore or one of coke? (b) How many megagrams of iron can
you make with the materials you have?
1
Expert's answer
2020-02-17T07:10:16-0500

The given reaction can be written as following:

Fe2O3 + 3C = 2Fe + 2CO

(a) To answer the first question, the limiting reagent should be determined. For the purpose, the number of moles of each reactant must be estimated:

n(Fe2O3) = m / Mr = 20.5 × 106 g / 160 g/mol = 1.23 × 105 mol

n(C) = m / Mr = 2.84 × 106 g / 3 × 12 g/mol = 7.9 × 104 mol

As carbon is a limiting reagent, more coke should be ordered.


(b) As carbon is a limiting reagent, 7.9 × 104 mol of iron will be produced. As a result:

m(Fe) = n(Fe) × 2 × Mr (Fe) = 7.9 × 104 mol × 2 × 56 g/mol = 8.85 × 106 g = 8.85 megagrams


Answer: (a) coke should be ordered, (b) 8.85 megagrams